[Hallc_running] [New Logentry] RC report / no meeting today / changing RC

mboer at jlab.org mboer at jlab.org
Sun Jul 28 13:50:08 EDT 2019

Logentry Text:
We are currently taking data at VCS kinematic I b. So far, we collected about 2.5 C at this kinematic and have 3 C of good data at matching VCS kinematic I a. 
Attached are results from our online replay scripts: missing mass spectra for all runs at kin I a and for the already replayed runs at kin I b. We are still improving some cuts such as hodo timing, trigger,..., and calorimeter calibration (at least) needs to be done. -> these plots give a lowest estimate of VCS signal candidates. 

Regarding running: we are running very efficiently with good beam, almost no down time, collecting data smoothly at 21 uA. We only had a false alert yesterday afternoon with LN2 level in HMS Q3 raising up above the alarm limit, and keeping going up (peak at 82%). However, after ~30 min of raising the level went down to values between alarm limits.

There is no RC meeting today. Mark Jones is taking over as RC right now. I would like to thanks all people who signed up at last minute for shifts when I was desperately begging last week, and in particular Hamlet M. for taking most day TO and Carlos Yero for taking most owl SL.  


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3711312
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190728/afc60750/attachment.html>

More information about the Hallc_running mailing list